the third term of a geometric progression is 4
product of first five term is?

Grade:11

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

\\t_3=4 \\=>ar^2=4 \\\prod_1^5t_i=a^4*r^{10}=4^5
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty